# MoA: Mixture of Sparse Attention for Automatic Large Language Model Compression

### v2

@@ 4.2 Automatic Optimization
MoA automatically selects the optimal elastic rule for each attention head to minimize accuracy losses across various sequence lengths under density budgets. Based on the profiling results, MoA first identifies Pareto front compression plans where any improvement in accuracy loss at one profile length would worsen another. To ensure the best generalization to lengths beyond those profiled, MoA then selects the plan that yields the minimum loss at an unseen length among the Pareto front solutions as the final plan.
Specifically, we utilize multi-objective optimization to search for a set of Pareto optimal compression plans across the profiled lengths. The objective for each length is to minimize the total accuracy loss while conforming to any user-defined density constraints. The objective is formulated as follows:
@@
Here, superscript $(N)$ denotes values at different lengths; $\mathbb{N}_{\text{profile}}$ and $\mathbb{N}_{\text{constr}}$ denote the sets of lengths for profiling and those subject to density constraints, respectively; $\mathbb{R}$ denotes the set of candidate rules; $\Delta L^{(N_{i})}$ denotes the accuracy loss due to compression; $d_{r_{h}}^{(N_{i})}$ denotes the density of rule $r_{h}$ at head $h$; $d_{\text{constr}}^{(N_{i})}$ denotes the average density constraint; $H$ denotes the total number of attention heads.
Such formulation corresponds to the classic multi-objective mixed-integer-planing problem, which can be effectively solved within minutes using existing linear solvers, like Gurobi (Gurobi Optimization, LLC, 2023). The detailed problem formulation and solving strategies are discussed in Appendix D.3.
Among the Pareto optimal compression plans, we select the one with the minimum loss at the unseen validation length as the optimal solution. This approach allows us to avoid profiling at every possible length while increasing the likelihood that the plan will generalize effectively to unseen lengths.
Thanks to this automatic pipeline, we efficiently get the elastic rules tailored for each attention head. With the pipeline, MoA minimizes the accuracy loss caused by attention sparsification, while conforming to user-defined density constraints.
@@ 5 Dataset and Supervision
In this section, we highlight the overlooked importance of calibration dataset design and its supervision objective in LLM compression. Calibration datasets are essential for sensitivity analysis across various compression techniques, including weight pruning (Men et al., 2024; Lee et al., 2024; Liu et al., 2023b) and quantization (Lin et al., 2023; Xiao et al., 2024b; Li et al., 2024a; Kim et al., 2023). In this work, MoA profiles the attention influence on the calibration dataset, which is crucial for subsequent automatic optimization.
@@
Current Approach. General language modeling datasets, such as human-written text corpus RedPajama (Computer, 2023), are commonly used as the calibration dataset. These datasets, supervised by next-token-prediction on the entire corpus, primarily capture attention patterns coherent with immediately preceding tokens. However, they lack long context dependencies, failing to address the global attention crucial for tasks like long-range retrieval.
@@
Moreover, a notable misalignment exists between the model response and the human-written supervision. Consequently, it leads to inaccuracies when using human responses to compute attention values and gradients during profiling. For example, given the same question, a human might answer ‘Blue’, while the model could generate ‘The blue color’. Using the human answer for supervision, attention influence is inaccurately quantified based on probability shift for predicting ‘Blue’; this diverges from the objective of maintaining crucial attention for the original model prediction, ‘The’. These inconsistencies arise from various factors, including mismatched positions, tones, and synonyms.
@@
MoA’s Approach. MoA enhances the calibration dataset by integrating long-range dependencies and model alignment. Specifically, we utilize the long-contextual MultiNews dataset (Fabbri et al., 2019), which includes summaries heavily dependent on long-range content. The summaries are generated by the original dense model and serve as supervision. Compared to current approaches that adopt human responses as the reference to calculate the loss, using the responses generated by the original model as the supervision can facilitate accurate influence profiling, thus benefiting the compression results.
@@
Approach Comparison. We evaluate our design’s effectiveness by varying dataset choices, supervision types, and summary references, while standardizing data item count and length to 50 and 8k words, respectively. Additional setups and evaluations are in Appendices A and B.3.1.
@@
We show the importance of long-range dependencies by comparing the MoA compression plan generated with different datasets and supervisory methods. In Table 2, RedPajama (Computer, 2023) represents the general language modeling dataset, while MultiNews (Fabbri et al., 2019) highlights long-range contexts by aggregating multiple documents on a single incident. Additionally, each MultiNews item includes a human-written summary, providing even stronger long-range dependencies and better performance. Calculating loss on the summary of MultiNews leads to significantly better performance, with a 60% increase in retrieval accuracy and a 0.98 decrease in perplexity.
@@ 6 Experiment
6.1 Setups
We brief the experiment setups here, with more details in Appendix A.
Baselines. We compare MoA with state-of-the-art static and dynamic sparse attention methods, including StreamingLLM (Xiao et al., 2024c), InfLLM (Xiao et al., 2024a) and H2O (Zhang et al., 2023). We define the density of an LLM as the ratio of the average in-memory KV-Cache length to the sequence length during the sparse decode stage. Notably, in MoA and StreamingLLM, KV-Cache length equals the attention span during the sparse prefill stage. In contrast, H2O use dense prefill. Besides, H2O and InfLLM require additional computations to dynamically determine the KV-Cache.
Models and Benchmarks. We evaluate on vicuna-{7b, 13b}-v1.5-16k models (Chiang et al., 2023) from LMSys and Llama-3-{8b, 70b}-Instruct-262k models (AI, 2024) from Gradient AI. For long-context retrieval, we use LongEval (Li et al., 2023) to test key-value retrieval accuracy with 100 data items per length level. For long-context understanding, we use LV-Eval (Yuan et al., 2024) and LongBench (Bai et al., 2023), which include 11 and 13 sub-datasets, respectively. For coherence testing, we measure perplexity on four long datasets (Dasigi et al., 2021; Fabbri et al., 2019; Li & Roth, 2002; Hovy et al., 2001; Mohler et al., 2016) with diverse tasks. Unless otherwise specified, performance experiments are restricted to eight A100-80GB GPUs over a 24-hour period, with OOM (Out-Of-Memory) and OOT (Out-Of-Time) conditions noted. Efficiency experiments measure the decode throughput on a single A100-80GB GPU at maximum batch sizes of respective methods.
MoA Settings. We restrict the number of distinct rules to at most two per model layer to ensure inference-time efficiency. We profile MoA on MultiNews (Fabbri et al., 2019) with model summaries at 2k, 4k, and 8k lengths. The optimal compression plan is selected with the validation dataset at 12k. Each model uses the same plan across all benchmarks and lengths. The models are not fine-tuned.
6.2 Accuracy-Throughput Trade-off
Figure 4 shows that MoA advances the Pareto Front in context retrieval accuracy and decode throughput across varied densities. At the same densities, MoA notably enhances throughput by 1.6×1.6$\times$1.6 × to 18.1×18.1$\times$18.1 × compared to H2O and InfLLM, due to its efficient static attention design. The throughput even outperforms StreamingLLM thanks to our customized GPU kernel. Additionally, MoA achieves notably higher retrieval accuracies across all densities. We conduct extensive evaluations for MoA ’s performance and efficiency on various benchmarks across context lengths from 4k to 256k and model sizes ranging from 7B to 70B in subsequent sections.
6.3 Performance
Figure 4: Accuracy-throughput trade-offs of various attention methods at different densities, tested on Vicuna-7B with 8k input length using one A100-80GB GPU on the LongEval dataset.
@@ 6.3 Performance
Table 4: Comparative analysis of retrieval accuracy, LV-Eval scores, LongBench scores, and perplexity for various models with different attention methods. All sparse methods employ 50% density in decode stage. H2O uses dense prefill, while StreamingLLM, InfLLM and MoA use sparse prefill. InfLLM for 70B model is excluded due to OOT issues.
@@
MoA outperforms state-of-the-art sparse attention methods across various model sizes and benchmarks, achieving comparable performance to the original dense model at 50% density.
Long-Context Retrieval. As shown in Table 4, MoA demonstrates a maximum of 8% relative accuracy drop (calculated as $\max\{1-\text{Acc.}_{\text{MoA}}/\text{Acc.}_{\text{Original}}\}$ across three lengths and LLMs), significantly less than the 87%, 58% and 44% for StreamingLLM, InfLLM and H2O. On average, the relative accuracy drop for MoA is under 1%, much less than others at 51%, 41% and 20%, respectively. Figure 5(a) shows that MoA retains over 90% retrieval accuracy up to 60k lengths, equaling the dense model’s effective context length. Note that it is done within 8k profiling and 12k validation. In contrast, the effective context lengths for H2O, InfLLM, and StreamingLLM are only 8k, <4k, and <4k, respectively. Appendix B.1.2 shows that MoA extends its effective context to approximately 3.9×3.9$\times$3.9 × the average KV-Cache length.
Long-Context Understanding. As shown in Table 4, MoA minimizes the maximum relative performance drop in LV-Eval and LongBench benchmarks to only 5% and 3%, respectively—much lower than the 36% and 18% experienced by StreamingLLM. H2O and InfLLM show maximum relative drops of 9%-17% and 3%-5% with higher efficiency costs. Similar trends show in perplexity tests, where MoA maintains less than 1% relative perplexity increase, while others exhibit 4%-13% increases. This trend holds for other densities, as shown in Appendices B.1.1 and B.1.3. Figure 9 and Table 8 further details the score with different tasks. MoA achieves comprehensive performance comparable to the original dense model, as well as H2O that requires higher efficiency cost. In contrast, StreamingLLM and InfLLM display inconsistent performance: it sometimes surpasses the original model in some tasks, while suffering noticeable degradation in others.
Longer-Context Generalization. By compressing within 12k lengths, MoA effectively generalizes to lengths of 32k-256k, as shown in Figure 5(b). At the extended lengths, MoA outperforms both InfLLM and StreamingLLM by 1.93.3×1.9-3.3$\times$1.9 - 3.3 × in retrieval accuracy and 1.21.4×1.2-1.4$\times$1.2 - 1.4 × in LV-Eval scores, demonstrating comparable performance to the original dense model.
@@
Figure 5: Comparative analysis at extended sequence lengths with different attention methods using Llama3-8B model. All methods employ 50% density in both prefill and decode stages.
Ablation Study. We evaluate the performance impact of different sparse mask search spaces in Table 3. Starting with a basic uniform mask, we observe significant enhancements by sequentially introducing heterogeneity: layers first, then heads, and finally elastic rules.
@@ 6.4 Efficiency
Table 5: Runtime efficiency of different methods on Vicuna-7B and 13B models. Efficiency improvements of MoA are ablated with four factors. All sparse attention methods use 50% density. Decode throughput (tokens per second) evaluated at the maximum batch capacity of an A100-80GB GPU.
@@
MoA shows high runtime efficiency with a manageable one-time compression overhead.
Runtime Efficiency. Table 5 compares the runtime efficiency of MoA over various attention methods and LLM frameworks, with the ablation of efficiency improvements brought by each design factor of MoA. At 50% density, MoA boosts the decode throughput by 6.6$\times$ to 8.2$\times$ compared to FlashAttention2. It outperforms H2O and InfLLM with 1.2×1.2$\times$1.2 × to 4.0×4.0$\times$4.0 × decode throughput improvements. Even compared to the highly system-level optimized vLLM framework (Kwon et al., 2023), MoA still achieves a 1.7×1.7$\times$1.7 × to 1.9×1.9$\times$1.9 × throughput increase. MoA also reduces total GPU memory by 1.2×1.2$\times$1.2 × to 1.4×1.4$\times$1.4 ×, are detailed in Appendix B.2. This throughput gain results from four main factors: static-sized KV-Cache during generation(3.0$\approx$ 3.0$\times$ 3.0 ×); reduced attention computations due to sparsity (1.5$\approx$ 1.5$\times$ 1.5 ×); increased batch sizes enabled by smaller KV-Cache memory (1.4$\approx$ 1.4$\times$ 1.4 ×); and our CUDA-implemented GPU kernel for MoA heterogeneous attention (1.2$\approx$ 1.2$\times$ 1.2 ×).
@@
Compression Pipeline Efficiency. MoA completes the automatic compression pipeline for the Vicuna-7B and 13B models within two hours. For the larger Llama3-70B model, the process requires 8.5 hours of real-time and 34.7 hours of GPU time. Details for each compression step overhead are in Appendix B.2.2.
6.5 Rules Discovered by MoA
We investigate MoA’s elastic rules for each head. As shown in Figure 10, masks in the initial and middle layers exhibit high density, aligning with the conclusions from previous research on LLM’s intrinsic dimensions (Valeriani et al., 2023) and layer sensitivities (Yuan et al., 2023). Conversely, in the final layers, most heads require low density, while few need high density. Figure 11 shows that layers with lower average mask density typically display a wider range of densities among heads, confirming the need for heterogeneity within the same layer. Further details and quantified insights are presented in Appendix C.
7 Conclusion and Future Work
MoA automates the selection of heterogeneous elastic masks for each attention head and input length, significantly extending the effective context length of LLMs by 3.9×3.9$\times$3.9 ×. It enhances retrieval accuracy by 1.5×1.5$\times$1.5 × to 7.1×7.1$\times$7.1 × over uniform sparse attention method and increases throughput to over 6×6$\times$6 × at 50% average density, maintaining performance on par with dense models in rigorous benchmarks.
Limitations and Future Work. Under an extremely low-density budget, MoA fails to maintain good performance. Designing a dynamic MoA method has the potential to address this issue, which we leave for future work. Using non-linear elastic rules with bounded attention spans is also worth exploring. Additionally, MoA’s profiling method can be adapted to evaluate the influence of weights and other activations, facilitating other compression methods such as weight and activation quantization (Lin et al., 2023; Xiao et al., 2024b; Li et al., 2024b).
Acknowledgement
@@
This work was supported by National Natural Science Foundation of China (No. 62325405, 62104128, U19B2019, U21B2031, 61832007, 62204164), Tsinghua EE Xilinx AI Research Fund, and Beijing National Research Center for Information Science and Technology (BNRist).
Appendix A Detailed Experiment Setup
A.1 Main Setup

### v3

@@ 4.2 Automatic Optimization
MoA automatically selects the optimal elastic rule for each attention head to minimize accuracy losses across various sequence lengths under density budgets. Based on the profiling results, MoA first identifies Pareto-front configurations where any improvement in accuracy loss at one profiled length would worsen another. To ensure the best generalization to lengths beyond those profiled, MoA then selects the configuration that yields the minimum loss at an unseen length among the Pareto front solutions as the final configuration.
Specifically, we utilize multi-objective optimization to search for a set of Pareto optimal configurations across the profiled lengths. The objective for each length is to minimize the total accuracy loss while conforming to any user-defined density constraints. The objective is formulated as follows:
@@
Here, superscript $(N)$ denotes values at different lengths; $\mathbb{N}_{\text{profile}}$ and $\mathbb{N}_{\text{constr}}$ denote the sets of lengths for profiling and those subject to density constraints, respectively; $\mathbb{R}$ denotes the set of candidate rules; $\Delta L^{(N_{i})}$ denotes the accuracy loss due to applying attention mask; $d_{r_{h}}^{(N_{i})}$ denotes the density of rule $r_{h}$ at head $h$; $d_{\text{constr}}^{(N_{i})}$ denotes the average density constraint; $H$ denotes the total number of attention heads.
This formulation corresponds to the classic multi-objective mixed-integer-programming problem, which can be effectively solved within minutes using existing linear solvers, like Gurobi (gurobi). The detailed formulation and solving strategies are discussed in Appendix E.3.
Among the Pareto-optimal MoA configurations, we select the one with the minimum loss at the unseen validation length as the optimal solution. This avoids profiling at every possible length and increases the likelihood of effective generalization to unseen input lengths.
@@ 5 Dataset and Supervision
We emphasize the often-overlooked role of calibration dataset design and supervision objectives in LLM compression. Calibration datasets enable effective sensitivity analysis across compression methods like weight pruning (Men2024ShortGPT; lee2024cats) and quantization (awq; smooth_quant). In this work, MoA profiles the attention influence on the calibration dataset, which is crucial for subsequent automatic optimization.
@@
Current Approach. General language modeling datasets, such as the human-written text corpus RedPajama (together2023redpajama), are commonly used as calibration datasets. These datasets, supervised by next-token prediction on the entire corpus, primarily capture attention patterns coherent with immediately preceding tokens. However, they lack long context dependencies, failing to address the global attention crucial for long-context tasks.
@@
Moreover, a notable misalignment exists between the model response and human-written supervision. Consequently, this leads to inaccuracies when using human responses to compute attention values and gradients during profiling. For example, given the same question, a human might answer ‘Blue’, while the model could generate ‘The blue color’. When using the human answer for supervision, attention influence is inaccurately quantified based on the probability shift for predicting ‘Blue’; this diverges from the objective of maintaining crucial attention for the original model prediction, ‘The’. These inconsistencies arise from various factors, including mismatched positions, tones, and synonyms.
@@
MoA’s Approach. MoA enhances the calibration dataset by integrating long-range dependencies and model alignment. Specifically, we utilize the long-context MultiNews dataset (Fabbri2019MultiNews), which includes summaries that depend heavily on long-range content. The summaries are generated by the original dense model and serve as supervision. Compared to current approaches that adopt human responses as the reference to calculate the cross-entropy loss $L$, using the responses generated by the original model as the supervision facilitates accurate profiling, thus benefiting the MoA configuration search.
@@
Approach Comparison. We validate our design with varied dataset choices, supervision types, and summary references, while standardizing the data item count and length to 50 and 8k tokens, respectively. Detailed setups and evaluations are in Appendices B and C.3.1.
@@
We show the importance of long-range dependencies by comparing the MoA configuration generated with different datasets and supervisory methods. In Table 2, RedPajama (together2023redpajama) represents the general language modeling dataset, while MultiNews (Fabbri2019MultiNews) highlights long-range contexts by aggregating multiple documents on a single incident. MultiNews additionally provides human-written summaries, enhancing long-range dependencies. Using MultiNews summaries for loss calculation significantly improves retrieval accuracy by 60% and reduces perplexity by 0.98.
@@ 6 Experiment
Figure 4: Accuracy-throughput trade-offs of seven attention methods at different densities, tested on Vicuna-7B with 8k input length using one A100-80GB GPU on the LongEval dataset.
@@
6.1 Setups
We briefly summarize the experimental setup here; more details appear in Appendix B.
Baselines. We compare MoA with state-of-the-art static and dynamic efficient attention methods, including StreamingLLM (xiao2023streamingLLM), InfLLM (xiao2024infllm) and H2O (Zhang2023H2O). We define the density as the ratio of the average in-memory KV-Cache length to sequence length during decoding. Notably, MoA and StreamingLLM use efficient prefill, while H2O and InfLLM use the original dense prefill, with additional computations to dynamically determine the KV-Cache for decoding.
Models and Benchmarks. We use vicuna-{7b, 13b}-v1.5-16k models (vicuna2023) from LMSys and Llama-3-{8b, 70b}-Instruct-262k models (meta2024llama3) from Gradient AI. We test long-context retrieval, understanding, and coherence abilities, using LongEval (lmsys2023longeval) retrieval accuracy, LV-Eval (yuan2024lveval), and LongBench (bai2023longbench) scores, as well as average perplexity across four datasets (Dasigi2021Qasper; Fabbri2019MultiNews; li-roth-2002-trec1; hovy-etal-2001-trec2; lcc). Performance experiments are restricted to eight A100-80GB GPUs over a 24-hour period, with Out-Of-Memory (OOM) and Out-Of-Time (OOT) conditions noted.
MoA Settings. We restrict the number of distinct rules to at most two per model layer to ensure inference-time efficiency. We profile MoA on MultiNews (Fabbri2019MultiNews) with model summaries at 2k, 4k, and 8k lengths. The optimal MoA configuration is selected using the validation dataset at 12k. Each model uses the same plan across all benchmarks and lengths. The models are not fine-tuned.
6.2 Accuracy-Throughput Trade-off
[…]
6.3 Performance
@@
Table 4: Comparative analysis of retrieval accuracy, LV-Eval scores, LongBench scores, and perplexity for various models with different attention methods. All methods employ 50% density in decode stage. H2O uses dense prefill; StreamingLLM, InfLLM and MoA use sparse prefill. InfLLM for 70B model is excluded due to OOT issues.
MoA outperforms state-of-the-art efficient attention methods across various model sizes and benchmarks, achieving comparable performance to full-attention model at 50% density.
Long-Context Retrieval. As shown in Table 4, MoA demonstrates a maximum of 8% relative accuracy drop (calculated as $\max\{1-\text{Acc.}_{\text{MoA}}/\text{Acc.}_{\text{Original}}\}$ across three lengths and LLMs), significantly lower than StreamingLLM (87%), InfLLM (58%), and H2O (44%). On average, MoA ’s relative accuracy drop is below 1%, substantially better than the other methods (51%, 41%, and 20%). Figure 5(a) shows that MoA retains over 90% retrieval accuracy up to 60k lengths, equaling the dense model’s effective context length. Note that it is done within 8k profiling and 12k validation. In contrast, the effective context lengths for H2O, InfLLM, and StreamingLLM are only 8k, <4k, and <4k, respectively. Appendix C.1.2 shows that MoA extends its effective context to approximately $3.9\times$ the average KV-Cache length.
Long-Context Understanding. Table 4 indicates that MoA reduces the maximum relative performance drop in LV-Eval and LongBench benchmarks to 5% and 3%, significantly less than StreamingLLM’s 36% and 18%. H2O and InfLLM incur maximum relative drops of 9%-17% and 3%-5%, respectively, but with greater efficiency costs. Figure 10 and Table 8 further show MoA’s consistent performance across subtasks, in contrast to the inconsistent performance of the baselines. In perplexity tests, MoA maintains less than a 1% increase in relative perplexity, whereas other methods see increases of 4%-13%. These trends persist at other densities (Appendices C.1.1, C.1.3).
[…]
@@
Figure 5: Experiments at extended sequence lengths with different attention methods using Llama3-8B model. All methods employ 50% density in both prefill and decode stages.
Ablation Study. We evaluate the performance impact of different sliding-window attention search spaces in Table 3. Starting with a basic uniform sliding-window method, we observe significant enhancements by sequentially introducing heterogeneity: layers first, then heads, and finally elastic rules.
@@ 6.4 Efficiency
Table 5: Decode throughput (tokens/s) of Vicuna-7B,13B, evaluated at the maximum batch capacity of an A100-80GB GPU. All efficient attention methods use 50% density. Efficiency improvements of MoA are ablated with four factors.
@@
MoA shows high runtime efficiency with a manageable one-time search overhead.
[…]
@@
Search Pipeline. MoA completes the automatic configuration search for the Vicuna-7B and 13B models within two hours. For the larger Llama3-70B model, the process requires 8.5 hours of wall time and 34.7 hours of GPU time. See Appendix C.2.4 for more details.
7 Conclusion
[…]
Acknowledgments
@@
This work was supported by the National Natural Science Foundation of China (No. 62325405, 62104128, U19B2019, U21B2031, 61832007, 62204164), the Tsinghua EE Xilinx AI Research Fund, and the Beijing National Research Center for Information Science and Technology (BNRist). We thank Yi Ge for valuable discussions.
Appendix A Notations and Definitions
To clarify the terminology and concepts introduced in our approach, we present the key notations and definitions used throughout the paper.
A.1 Attention Span
The attention span of each attention head is defined as the sum of a sliding-window length and a fixed prefix of always-visible tokens (attention sink). In the attention mask:
  • Sliding Window: A contiguous diagonal region indicating recent tokens each head can attend to.
  • Prefix: The first 64 tokens of the input sequence, which remain visible to all attention heads regardless of the sliding-window length.
A.2 Elastic Rule
An elastic rule is a function $f(\cdot)$ parameterized by $\alpha$ and $\beta$, specifying how the attention span scales with the input sequence length $N$ for a single head. Formally, the attention span $\mathcal{S}$ given by an elastic rule is:
$\mathcal{S}(N;\alpha,\beta)=\alpha\times N+\beta$ (6)
Different attention heads employ different elastic rules by adopting different sets of $\alpha$ and $\beta$, resulting in heterogeneous attention spans optimized per head.
A.3 MoA Configuration
An MoA configuration is the collection of elastic rules of all attention heads of all layers in a model.
A.4 Attention Mask
A.4.1 Visualization
An attention mask is a binary mask applied to each head individually to gate its attention computation. It is defined by the elastic rule and indicates which tokens are visible (unmasked) or pruned (masked).
In Figure 3, each attention mask has three visual elements:
Visual Element Meaning Contribution to Span
Vertical purple stripe Initial unmasked tokens (prefix) 64 tokens (fixed)
Diagonal purple band Sliding window region Window length
White cells Pruned tokens 0
As shown in the figure, each grid represents a candidate attention mask:
  • Vertically aligned masks represent masks generated from different elastic rules at the same input length.
  • Horizontally aligned masks represent masks generated from the same elastic rule at different input lengths.
A.4.2 Application of Attention Masks
Attention masks for all heads are applied in parallel—each mask independently gates attention computation and the corresponding key-value (KV) cache entries. There is no overlap or sequential application between masks across different heads.
A.5 Correlations Between Notions
The following list summarizes the correlations between notions. Specifically:
  1. 1.
    MoA configuration and elastic rule: An MoA configuration is the elastic rules of all attention heads of all layers in an LLM.
  2. 2.
    Elastic rule and attention Span: Given an elastic rule parameterized by $(\alpha,\beta)$ and input length $N$, the attention span is computed using Equation 6.
  3. 3.
    Attention span and attention mask: The computed span directly determines the shape and extent of the sliding-window portion of the attention mask.
Appendix B Detailed Experiment Setup
B.1 Main Setup